A 66 kg adult sits at one end of a 8.0 m long board. His 25 kg child sits on the other end.
(a) Where should the pivot be placed so that the board is balanced, ignoring the board's mass?

____ m (measured from the adult)
(b) Find the pivot point if the board is uniform and has a mass of 15 kg.

_____m (measured from the adult)

Answer by stanbon(75887)

A 66 kg adult sits at one end of a 8.0 m long board. His 25 kg child sits on the other end.
(a) Where should the pivot be placed so that the board is balanced, ignoring the board's mass?
Equation::
66/x = 25/(8-x)
66*8-66x = 25x
66*9 = 91x
x = 6.53 meters from the 25 kg child
-----------------------------
1.47m (measured from the adult)
(b) Find the pivot point if the board is uniform and has a mass of 15 kg.
1.47m (measured from the adult)
------------
Cheers,
Stan H.
-----------